The length of a rectangle is 5 m greater than the width. The area is 300 sq m. Find the dimensions.
Answer by checkley79(3341)

L=W+5
A=LW
300=(W+5)W
300=W^2+5W
W^2+5W-300=0
(W-15)(W+20)=0
W-15=0
W=15 AND.
L=15+5
L=20 ANS.
PROOF:
300=15*20
300=300